OpenStudy (anonymous):

What percent of 60 is 12? Show your work.

OpenStudy (anonymous):

just divide 12 by 60 \[\frac{12}{60}=?\] You will get a decimal value but then just turn the decimal into percent

OpenStudy (anonymous):

let number be x then 60 percent of x=12 x=(12*100)/60=20 ans is 20

OpenStudy (anonymous):

12/60 *100=20

OpenStudy (anonymous):

12(100)/60=20